HUBBLE SPACE TELESCOPE DAILY REPORT #946 PERIOD COVERED: 0800 EST Daily Status Report as of 221/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WFC 4292 (Determination of the Masses and Densities of Pluto and Charon - Cycle 3) The Wide Field Camera was used to make a total of six observations of the planet Pluto, its moon Charon, and a nearby reference standard star. These images will be used to determine the motion of Pluto about the Pluto/Charon barycenter which will make it possible to compute the mass ratio of the two bodies and put constraints on their masses and densities. All six observations were received and displayed. No anomalies were reported, and the images were reported to look normal. 1.2 Completed Three Sets of HSP/POL 1097 (X-Ray Binaries Part 2 - Cycle 2 High) The High Speed Photometer was used to make a series of twenty photometric observations of the X-Ray binary star Cygnus X-1 (aka, HDE226868) taken with several different filters. All of the observations were received and displayed in the OSS. They all were reported to look good. 1.3 Completed WFC 4848 (GTO Low Latitude Parallel Program - Cycle 3 High) The Wide Field Camera was used to make two high latitude medium-deep sky survey observations taken in parallel with the prime science observations. Both observations were received and displayed. They were reported to look okay. 1.4 Completed PC 4712 (WF/PC Cycle 2 Calibration: Measles Monitor) The Planetary Camera was used to make three Earth Calibration Flat images to monitor fo a form of contamination known as measles. The observations used filter F517N, both with and without filter F122M. The OSS monitored the observation and assessed the data quality. The first image was reported to be saturated, and the last two were reported to look normal. 1.5 Completed FOS/BL 4976 (UV Rotational Light Curves for Pluto and Charon's UV Spectrum - Split) The Faint object Spectrograph was used to make a TALED exposure subsequent to other spectral observations of the planet Pluto and its moon Charon. The TALED observation was received in real time. It was reported to look okay. After a corrective SIP maneuver was uplinked to the spacecraft, spectral observations were made of the planet Pluto. The observations were executed in fine lock. All of the observations were reported to look nominal. 1.6 Completed FOS/RD 4123 (FOS Spectral Flat Fields Monitor, Revised) The Faint Object Spectrograph (Red Side Only) was used to make standard spectral flat calibration observations using standard star WD0501+527 to track the time dependence of the spectral response in the UV. Two status buffer error messages were generated during the first observation. These referred to the failure, and then the success, of a "speed check". This problems did not appear to affect the observations. The three spectral observations were reported to look normal. 1.7 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.8 Completed GHRS/2 3845 (Deciphering the UV Emission Lines in the T Tauri systems) The High Resolution Spectrograph was used to make six ultraviolet spectral observations of the T-Tauri class star DF-Tauri to study the width and shape of its emission lines which will put constraints on the type of mechanism producing these lines. One wavelength calibrations and six spectral observations were received and displayed. No problems were reported. 1.9 Completed HSP/UV1 1091 (Ultraviolet Pulsations From X-Ray Pulsars) The High Speed Photometer was used to make single color photometric observations of the pulsating X-Ray (T-Tauri Class) binary star known as 4U1145-619. This is to find out if this object give off ultraviolet, as well as X-Ray pulsations. The observations were executed as scheduled, and no problems were reported. 1.10 Completed FGS 5055 (A Critical Test of the Galactic Escape Velocity at R(SUN): Cycle 1 Obs.) The Fine Guidance Sensors were used to make astrometric observations of the high velocity star G166-37 to determine its galactic escape velocity. The FGSs will be used in the POS mode to measure the target's trigonometric parallax and proper motion. The astrometric observations were monitored in the OSS, and no problems were seen. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 19 Successful : 19 Scheduled Reacquisitions : 22 Successful : 22 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 39 Successful : 38 U/D 1,2 Role Delay failed all three iterations of tracker #1 at 221/0650Z. Full Maneuver which followed was okay. 2.3 Significant HSTARs: None 2.4 Operations Note: An attempt to remove the uniform layer of contamination from the WF/PC CCD windows, which have been accumulating over the last year, was performed by executing a real time Flash Decontamination Procedure. The twenty step procedure was carried out in four uplink windows between 214/1818Z and 215/0026Z. The WF/PC CCD, Hot Junction and Radiator temperatures were monitored throughout (as down links permitted) and compared against the anticipated thermal performance to decide which if any pre-planned contingency paths were to be taken (none were) and how the command timing should be spaced within the allowable tolerances. As noted in an earlier memo, the WF/PC M1 mirror heated was turned off approximately seven hours before starting this procedure to minimize the temperature gradient across the radiator. This, as expected, resulted in reducing the dispersion in the CCD temperatures during the decontamination period. After warming up, the heat pipe heaters were commanded off at the opening of the uplink window in which that activity was planned. It was anticipated that the temperature of the P7 CCD (used to track the warm-up performance) would be near it's target temperature of -2C at the opening of that window. It was seen to be somewhat warmer, at +1.3c, and hence the heat pipe heaters were turned off by ground command. Through out the one hour forty minutes of the decontamination period the CCD temperatures, with the exception of W1 remained at about 3C +/- 2 Degrees. This met the goal of this procedure. However, unexpectedly, W1 ran approximately 6 to 8 degrees colder than the other CCDs. The reason for this , at this time is not clear. Nor is it known how effective the decon attempt will be for W1. Internal flats will be taken to assess how well the procedures worked. At the start of the cool-down, when the TEC power was reapplied, the TEC voltages were below their normal operating voltage by about one volt( at about 18 volts). The TEC voltages vary with the radiator temperature. Since the M1 heater had been off for about 11 hours, the radiator at that time had fallen, as expected to approx. -49C. It is believed that the low voltage condition was due to the large thermal gradient across the CCD junctions. The TECs cooled down as expected, and within a few minutes after turn-on the TEC voltage increased to the nominal operational range. MOSES has opened an HSTAR on this low-voltage condition and SESD will assist in quantifying this condition. It is likely that the limit set used for the TEC voltages, with the M1 mirror heater off, may not be appropriate during TEC transitional periods. At this time the CCDs have cooled and are nearly to their regulation temperatures. Note that the M1 mirror heater has been intentionally left powered off. If SIB's analysis of the internal flats shows that the decontamination was successful then the M1 mirror heater will be turned back on. If a repeat of the decontamination is needed the M1 mirror heater will already be powered off at the start. /GS/STScI 8/6/93 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Three Sets of WFC 4713 (WF/PC Cycle 2 Calibration: WFC Internal Delta Flats) 3.2 Complete Two Sets of FGS 4283 (Mass Ratio Determination For Doubles With Substellar Mass Members) 3.3 Complete WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) 3.4 Complete PC 4929 (WF/PC Cycle 2: Improved Dark Calibration) 3.5 Complete GHRS/2 5069 (Empirical Determination of the Wind Velocity and Density Laws) 3.6 Complete PC/WFC 4109 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor) 3.7 Complete FGS 4983 (Orbital Motions of Close Binary Stars - Cycle 2 High) 3.8 Complete PC 4714 (WF/PC Cycle 2 Calibration: WFC Internal Delta Flats) 3.9 Complete WFC 5080 (A Study of the Arcsecond Gravitational Lens 1422+321 - Cycle 3 High) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-